OpenStudy (anonymous):

The points (5,2), (-10,-2), and (t,01) lie on the same line. Find the value of t.

OpenStudy (anonymous):

please help me.....

OpenStudy (tamtoan):

if they are on the same line, they have the same slope...and from first 2 points you can find the slope, then use first point and third point ..slope will have t in it...then solve for t.

OpenStudy (anonymous):

oh im sorry the question was: The points (5,2), (-10,-2), and (t,-1) lie on the same line. Find the value of t. not 01.

OpenStudy (tamtoan):

just do what i guide you there, you should get the answer

OpenStudy (anonymous):

How do i find the slope again?

OpenStudy (tamtoan):

formule of finding slope between 2 points (x1, y1) and (x2,y2) is : m = (y2-y1)/(x2-x1) there.

OpenStudy (tamtoan):

formula

OpenStudy (anonymous):

I keep getting a wrong answer. Can u tell me the correct answer so i see where i went wrong? thanks.

OpenStudy (tamtoan):

show me your work and i'll help you find out where you get it wrong.

OpenStudy (anonymous):

Sorry, my interent isnt working weel. My outcome of the problem was a slope of 4/5.Could i watch u work it out once.

OpenStudy (tamtoan):

can you just show me your work finding the slope with the first 2 points on here..use draw to write it..it's much easier and faster.

OpenStudy (anonymous):



OpenStudy (tamtoan):

yeah you got that correct , you can cancel out the negative signs since they both have on top and bottom.. now use the same approach, what is the slope for first point and third point?

OpenStudy (anonymous):

I keep getting stuck. Do I plug in the numbers i got for the last equation of do i do this: -1-1/t-10= ? IHow do i do this with the t?

OpenStudy (anonymous):

please, can I have the answer?

OpenStudy (tamtoan):

the last point, your x coordinate is t and your y coordinate is -1

OpenStudy (tamtoan):

let me write down the points again so i don't make mistake : first point (5,2) second point (-10,-2) third point (t, -1) you already got slope m = 4/15 use first point and third point you also have m = (-1 - 2)/(t-5) = -3/(t-5) since they are on the same line..so the slope is the same: so you have 4/15 = -3/(t-5) can you take it from here and solve for t ?

OpenStudy (anonymous):

i cant remember how to clear the t.

OpenStudy (tamtoan):

multiply accross ...you will have 4(t-5) = -3x15 so 4t-20 = -45 now it's easy right? :)

OpenStudy (anonymous):

Okay...so the answer to this problem is t=-6.25?

OpenStudy (anonymous):

But don't I need a (0,0) anseer?

OpenStudy (tamtoan):

your t = -25/4 ...that's correct ...what do you mean by a (0,0) answer ? the question ask for the value of t , right?

OpenStudy (anonymous):

so lets say i got t=-7 for another answer....do i put it as -7/1 for the answer?